Given the following pairs of z-values, find the area under the normal curve between each pair of z-values. Refer to the table in Appendix B.1. (Round the final answers to 4 decimal places.)
a. z = -1.95 and z = 0.75
b. z = -1.05 and z = -0.75
c. z = 1.77 and z = 2.98
d. z = -2.43 and z = 1.43
Answer #1

Solution:

(a)  z = -1.95 and z = 0.75

P(-1.95<Z<0.75)=P(Z<0.75)-P(Z<-1.95)

P(-1.95<Z<0.75)=0.7734 - 0.0256=0.7478

P(-1.95<Z<0.75)=0.7478

(Round to four decimal places)

(b) z = -1.05 and z = -0.75

P(-1.05<Z<-0.75)=P(Z<-0.75)-P(Z<-1.05)

P(-1.05<Z<-0.75)=0.2266-0.1469=0.0797

P(-1.05<Z<-0.75)=0.0797

(Round to four decimal places)

(c) z = 1.77 and z = 2.98

P(1.77<Z<2.98)=P(Z<2.98)-P(Z<1.77)

P(1.77<Z<2.98)=0.9986-0.9616=0.0370

P(1.77<Z<2.98)=0.0370

(Round to four decimal places)

(d)  z = -2.43 and z = 1.43

P(-2.43<Z<1.43)=P(Z<1.43)-P(Z<-2.43)

P(-2.43<Z<1.43)=0.9236-0.0075=0.9161

P(-2.43<Z<1.43)=0.9161

(Round to four decimal places)

PLEASE RATING THE ANSWER. THANK YOU.....!!!!